OpenStudy (anonymous):

our daily workout plan involves a total of 60 minutes of jogging and biking. You burn 10 calories per minute jogging and 5 calories per minute biking. How many calories will you burn in your 60−minute workout if you jog for 30 minutes?

OpenStudy (anonymous):

30 *10 = 300 calories 30*5 = 150 calories Total 300 +150 = 450 calories